# Navigating Public Sentiment in the Circular Economy through Topic Modelling and Hyperparameter Optimisation

### v1

@@ 2.4 Evaluation Metrics
NPMI normalises PMI to range between -1 and 1, where 1 indicates perfect coherence, 0 indicates independence, and -1 indicates perfect incoherence.
@@
We then take the average of the scores of all bigrams within a topic as the coherence score for that topic. Finally, we average the coherence scores across all topics to obtain the overall coherence score for each model.
@@ 2.5 Circular Economy
One strand of research examined the public recognition and understanding of the concept of the CE from broader social perspective [45, 46]. Some authors focused on exploring consumers’ attitudes and purchase intentions towards the CE, including preferences for remanufactured products and circularly packaged products [47, 48]. Furthermore, the literature analysed companies’ perspectives on the transition to a circular economy, how they incorporate circular economy initiatives into their production operations, and the obstacles they face in this process [49, 50, 51]. To ensure broader coverage of CE-related news data, 53 keywords were included in the experiment to ensure Guardian news data can be obtained more widely.
Digitalisation has significantly reshaped firm’s production models in recent years [52], extending its influence beyond information technology to various aspects, including the circular practices. For instance, digital technologies play a crucial role in achieving carbon neutrality by enhancing product value and reducing pollution [53]. The application of digital technology not only improves the efficiency of the CE but also allows companies to broaden the scope of its applications and effectively utilise scarce resources [54]. Furthermore, integrating the circular economy concept into product and service design substantially enhances consumption potential by decoupling growth from resource availability [48]. Previous research has identified a positive relationship between digitalisation and the development of the CE [54]. Tools such as big data analytics and the Internet of Things (IoT), when combined with digital technology, enable monitoring and tracking of product life cycle data, thus minimising waste and enhancing recycling rates [55]. Artificial intelligence enhances productivity and promotes CE development through optimised processes and real-time data analysis [54]. Additionally, machine learning anticipates uncertainties in the design and production process, identifying deficiencies in the CE system through real-time monitoring [56]. Therefore, analysing the circular economy from a digital perspective can assist companies in achieving a low-carbon transformation and effectively capturing public views and attitudes towards the circular economy.
@@ 3.1 Framework
Figure 2 describes the proposed framework for the interdisciplinary study, with a particular emphasis on investigating trends and concerns within public opinion regarding the CE. Firstly, we collected textual posts from Twitter, Reddit, and The Guardian platforms to reflect the opinions and attitudes of three distinct strata of the public. Then, it is worth to note that to ensure compliance with data regulations, we utilised official APIs to retrieve the relevant data. Next, we conducted exploratory data analysis with the purpose of gaining a preliminary understanding and getting to know the data landscape before taking necessary data preprocessing, i.e., data cleaning. In Figure 2, the red dashed rectangle contains the processes related to data collection, which will be discussed in Section 3.2. Meanwhile, the green dashed rectangle includes hyperparameter optimisation for topic modelling, which will be discussed in Section 3.3.
Figure 2: The proposed multidisciplinary framework for using hyperparameter optimisation on topic modelling in the context of circular economy.
@@ 3.2 Datasets
CE is a broad topic that involves different groups of people, such as the public, professionals, and officials. Analysing the various opinions and interests in CE across different groups requires us to collect data from a wide spectrum of society. To achieve this goal, we collected textual posts from Twitter and Reddit, respectively representing the opinions of the public and professionals. Meanwhile, we also scraped The Guardian news to analyse official opinions on CE.
The process of data scraping for this research can be summarised as the following steps:
@@
    Search suitable websites containing data from various segments of the public individually. Ensure a comprehensive understanding of diverse perspectives.
@@
    Check the policy of the selected websites to ensure that data scraping is permitted for research purposes and that the process complies with their regulations.
@@
    Follow the university’s rules to submit ethic approval application.
@@
    Investigate the usage of the APIs offered by websites. Data scraping can be challenging, especially for websites that reserve large amounts of data, as searching for user-customised data is time-consuming. Additionally, different websites may design different APIs with varying arguments (e.g., keywords, time span) to support efficient data scraping. Therefore, it is important to be familiar with the API usage to ensure that the downloaded data meets expectations.
In machine learning, data preprocessing is a critical step for models to achieve satisfactory performance. This step is also vital for topic modelling because we have found that the quality of the original scraped data is poor and not suitable for direct use in topic models. For example, website links commonly appear in social media posts and news; however, they are less informative for topic modelling and may adversely affect model performance. In this research, data preprocessing primarily involves several targets:
  1. 1.
    Stopwords removal
  2. 2.
    Punctuation removal
  3. 3.
    Email and URL removal
  4. 4.
    Special symbols removal
  5. 5.
    Lemmatisation
  6. 6.
    Check whether the text length exceeds the model limit
@@ 3.3 Hyperparameter Optimisation for Topic Modelling
Hyperparameters in machine learning denote the parameters that need to be set before the model is trained or fitted. In other words, a machine learning algorithm can be instantiated by configuring it with a set of hyperparameter values. These values control the behaviour of algorithms and also affect aspects such as model complexity and convergence. Therefore, different sets of hyperparameter values can produce different results.
Hyperparameter selection inherently is an iterative process of trial and error, and the selection usually relies on a quantitative metric. In our research, to compare different sets of hyperparameter values, we used the coherence (NPMI) value, which ranges from -1 to 1; a value of 1 indicates the best coherence. Manual tuning of hyperparameters is very time-consuming and relies on machine learning intuition and experience. Therefore, many advanced approaches have been proposed to automate the process of hyperparameter optimisation, such as CMA-ES [57] and TPE [58]. However, these approaches usually perform well and exhibit their strengths on relatively large hyperparameter search spaces.

### v2

@@ 2.4 Evaluation Metrics
NPMI normalises PMI to range between -1 and 1, where 1 indicates perfect coherence, 0 indicates independence, and -1 indicates perfect incoherence. We then take the average of the scores of all bigrams within a topic as the coherence score for that topic. Finally, we average the coherence scores across all topics to obtain the overall coherence score for each model.
To further enhance the measurement of topic coherence, Röder et al. [16] introduced the $C_{\text{NPMI}}$ metric, which extends NPMI by averaging its overall word pairs in a topic’s top $N$ words. The $C_{\text{NPMI}}$ coherence score for a single topic is calculated as:
$C_{\text{NPMI}}=\frac{1}{\binom{N}{2}}\sum_{i<j}\text{NPMI}(\omega_{i},\omega_{j}),$ (13)
where $\omega_{i}$ and $\omega_{j}$ are the representative words of the topic. This metric effectively captures the overall semantic coherence by considering the pairwise associations between all top terms within a topic. The final coherence score for the model is then obtained by averaging the $C_{\text{NPMI}}$ scores across all topics.
By employing $C_{\text{NPMI}}$, we obtain a more robust evaluation of topic quality, as it mitigates the biases inherent in PMI and NPMI when applied to individual word pairs. This comprehensive approach to measuring coherence has been widely adopted in recent topic modelling research due to its effectiveness in reflecting human judgments of topic interpretability.
Diversity
Topic diversity is a crucial metric for evaluating the distinctiveness and breadth of topics produced by a topic model. It assesses the extent to which topics are unique and non-overlapping, ensuring that the model captures a wide range of topics present in the corpus [46]. To quantify diversity, we utilise the Inverted Rank-Biased Overlap (Inverted RBO) metric [47], which measures the dissimilarity between ranked lists of words representing different topics.
Rank-Biased overlap (RBO) is a similarity measure for ranked lists that accounts for top-weightedness, meaning higher-ranked items contribute more to the overall score [47]. Given two topics $t_{i}$ and $t_{j}$, each represented by an ordered list of their top $N$ words, the RBO is defined as:
$\text{RBO}(t_{i},t_{j})=(1-p)\sum_{k=1}^{N}p^{k-1}\cdot\frac{|W_{t_{i}}^{(k)}\cap W_{t_{j}}^{(k)}|}{k},$ (14)
where $W_{t}^{(k)}$ is the set of the top $k$ words in topic $t$. $p$ is the persistence parameter $(0<p<1)$, controlling the weight decay; a higher $p$ gives more weight to lower-ranked items. To measure diversity, we compute the Inverted RBO between all pairs of topics, which reflects the dissimilarity:
Inverted $\text{Inverted\ RBO}(t_{i},t_{j})=1-\text{RBO}(t_{i},t_{j}).$ (15)
The overall diversity $D$ of the topic model is then calculated as the average Inverted RBO across all unique topic pairs:
$D=\frac{2}{|T|(|T|-1)}\sum_{i=1}^{|T|}\sum_{j=i+1}^{|T|}\left(1-\text{RBO}(t_{i},t_{j})\right),$ (16)
where $|T|$ is the total number of topics. A higher $D$ value indicates greater diversity among topics, signifying that the model captures a wider array of distinct topics with minimal redundancy.
Perplexity
This is a standard metric used to evaluate the predictive performance of probabilistic models in natural language processing [11]. It gauges how well a model predicts a sample of data, with lower perplexity values indicating better generalisation to unseen data.
In the context of topic modelling, perplexity assesses the model’s ability to assign high probabilities to words in unseen documents. Formally, the perplexity $\mathcal{P}$ of a held-out test set is defined as:
$\mathcal{P}=\exp\left(-\frac{\sum_{d=1}^{D}\sum_{w=1}^{N_{d}}\log P(w_{d,w}\,|\,\theta_{d})}{\sum_{d=1}^{D}N_{d}}\right),$ (17)
where $D$ is the number of documents in the test set. $N_{d}$ is the number of words in document $d$. $w_{d,w}$ is the $w$-th word in document $d$. $P(w_{d,w}\,|\,\theta_{d})$ is the probability of word $w_{d,w}$ given the document’s topic distribution $\theta_{d}$.
To compute perplexity, we ensure numerical stability by avoiding zero probabilities, typically by applying a small constant $\epsilon$ (e.g., $1\times 10^{-10}$) to clip the probabilities:
$P(w_{d,w}\,|\,\theta_{d})=\max\left(P(w_{d,w}\,|\,\theta_{d}),\epsilon\right).$ (18)
This adjustment prevents the logarithm from becoming undefined due to zero probabilities. The perplexity metric effectively evaluates the exponential of the negative average log-likelihood per word, providing insight into the model’s predictive uncertainty [13]. A lower perplexity score implies that the model is more adept at predicting word occurrences in unseen documents [46].
Pareto Front
The Pareto front represents the set of optimal solutions in Multi-objective Optimisation where no objective can be improved without degrading another [48, 49]. In topic modelling, this concept helps us balance conflicting metrics such as coherence, diversity, and perplexity [46]. A solution $\mathbf{x}^{*}$ is Pareto optimal if there is no other solution $\mathbf{x}$ such that:
$\forall m,\quad f_{m}(\mathbf{x})\leq f_{m}(\mathbf{x}^{*})\quad\text{and}\quad\exists m,\quad f_{m}(\mathbf{x})<f_{m}(\mathbf{x}^{*}),$ (19)
where $f_{m}$ are the objective functions representing different evaluation metrics. By mapping these solutions, we identify hyperparameter settings that offer the best trade-offs, aiding in the selection of models that align with specific performance goals.
Tree-structured Parzen Estimator Application in Multi-objective Optimisation
The Tree-structured Parzen Estimator (TPE) is a Bayesian optimisation algorithm effective for hyperparameter tuning [50]. While originally designed for single-objective optimisation, TPE can be adapted for multi-objective problems by optimising a scalarised objective function [51]. In this research, we employ TPE to navigate the hyperparameter space efficiently, aiming to find models that balance coherence, diversity, and perplexity.
Although the Pareto front provides a set of non-dominated solutions, selecting a single optimal point requires further criteria. We utilise the Ideal Point Method [52], which defines an ideal point $\mathbf{z}^{*}$ composed of the optimal values for each objective:
$\mathbf{z}^{*}=\left(\min_{\mathbf{x}}f_{1}(\mathbf{x}),\min_{\mathbf{x}}f_{2}(\mathbf{x}),\ldots,\min_{\mathbf{x}}f_{M}(\mathbf{x})\right),$ (20)
where $f_{m}(\mathbf{x})$ are the objective functions. For each solution $mathbf{x}$ on the Pareto front, we calculate the Euclidean distance (schema shown in Figure 2) to the ideal point:
$d(\mathbf{x})=\sqrt{\sum_{m=1}^{M}\left(f_{m}(\mathbf{x})-z^{*}_{m}\right)^{2}}.$ (21)
We select the solution with the minimal distance $d(\mathbf{x})$ as the optimal hyperparameter configuration. This approach allows us to convert the multi-objective problem into a single-objective one, making it compatible with TPE optimisation [50, 51]. By integrating TPE with the Ideal Point Method, we effectively handle multiple objectives and identify hyperparameters that offer the best trade-offs among coherence, diversity, and perplexity.
Figure 2: Schematic representation of the Pareto front, the ideal point, and the final optimal solution minimising two contradictory objectives $f_{1}$ and $f_{2}$ [52].
@@ 2.5 Circular Economy
One strand of research examined the public recognition and understanding of the concept of the CE from a broader social perspective [53, 54]. Some authors focused on exploring consumers’ attitudes and purchase intentions towards the CE, including preferences for remanufactured products and circularly packaged products [55, 56]. Furthermore, the literature analysed companies’ perspectives on the transition to a circular economy, how they incorporate circular economy initiatives into their production operations, and the obstacles they face in this process [57, 58, 59]. To ensure broader coverage of CE-related news data, 53 keywords (Appendix Table 18) were included in the experiment to ensure Guardian news data can be obtained more widely.
Digitalisation has significantly reshaped a firm’s production models in recent years [60], extending its influence beyond information technology to various aspects, including circular practices. For instance, digital technologies play a crucial role in achieving carbon neutrality by enhancing product value and reducing pollution [61]. The application of digital technology not only improves the efficiency of the CE but also allows companies to broaden the scope of its applications and effectively utilise scarce resources [62]. Furthermore, integrating the circular economy concept into product and service design substantially enhances consumption potential by decoupling growth from resource availability [56]. Previous research has identified a positive relationship between digitalisation and the development of the CE [62]. Tools such as big data analytics and the Internet of Things (IoT), when combined with digital technology, enable monitoring and tracking of product life cycle data, thus minimising waste and enhancing recycling rates [63]. Artificial intelligence enhances productivity and promotes CE development through optimised processes and real-time data analysis [62]. Additionally, machine learning anticipates uncertainties in the design and production process, identifying deficiencies in the CE system through real-time monitoring [64]. Therefore, analysing the circular economy from a digital perspective can assist companies in achieving a low-carbon transformation and effectively capturing public views and attitudes towards the circular economy.
@@ 3.1 Framework
Figure 3: The proposed framework employs single-objective and multi-objective hyperparameter optimisation on topic modelling in the context of the circular economy. SOO and MOO respectively denote single-objective and multi-objective optimisation.
Figure 3 depicts the proposed framework for this interdisciplinary research, which employs topic modelling with twin hyperparameter optimisations and places particular emphasis on investigating public attention regarding the CE.
Firstly, we identified the research context, CE. To explore public attention regarding CE, we divided the sources into three groups: the general public, professionals, and official entities. Then, we collected textual posts from Twitter, Reddit, and The Guardian platforms to reflect the attention of these three distinct strata of the public. It is worth noting that, to ensure compliance with data regulations, we utilised official APIs to retrieve the relevant data. More details regarding the datasets can be found in Section 3.2.
Next, we data.Next,an we conducted exploratory data analysis to gain a preliminary understandisetdata landscape before performing necessary datad As shown in Figure 3, LDA, CorEx, and BERTopic represent the three distinct topic models used in this research, each chosen for its specific motivation and approach to topic modelling. These models were subjected to single-objective hyperparameter optimisation using grid search, allowing for a relatively thorough, simple, and systematic exploration of the hyperparameter space to identify the optimal values. Coherence was selected as the objective metric, as it is a well-known and commonly used measure of topic quality in topic modelling.
Many advanced approaches have been proposed to automate the hyperparameter optimisation process, such as CMA-ES [65] and TPE [66]. However, these approaches usually perform well and exhibit their strengths on relatively large hyperparameter search spaces. According to our designed hyperparameter search space, we consider grid search a promising option as it can exhaustively explore all solutions explicitly.
However, single-objective hyperparameter optimisation can potentially overlook other critical aspects, such as the diversity and perplexity of the generated topics. Therefore, we proposed a twin hyperparameter optimisation framework. We selected the best-performing topic modelling approach from the single-objective optimisation results to further apply multi-objective optimisation using the Tree-structured Parzen Estimator algorithm on a finer hyperparameter grid. We then analysed the two sets of topics generated from the models optimised by both single-objective optimisation (SOO) and multi-objective optimisation (MOO), with the latter serving as auxiliary information, offering an alternative perspective to support our investigation of CE. Additional details on SOO and MOO can be found in Section 3.3.
@@ 3.2 Datasets
The CE is a broad topic that involves various stakeholders, including the general public, professionals, and officials. Analysing the diverse opinions and interests in CE across these groups is essential for understanding public attention, which necessitates collecting data from a wide spectrum of society. To accomplish this, textual posts were collected from Twitter and Reddit, representing the perspectives of the general public and professionals, respectively. Additionally, news articles from The Guardian were gathered to analyse official opinions and attention regarding CE. We collected a total of 22,106 documents related to the CE and related keywords from three different sources, strictly following each platform’s API usage policy, and each dataset was based on a fixed time range of 1999 to 2023.
The process of data scraping for this research can be summarised in the following steps:
@@
    Search for suitable websites containing data related to CE from various segments of the public to form a comprehensive dataset, ensuring that subsequent analysis can cover a wide range of perspectives on CE.
@@
    Check the policies of the selected websites to ensure that data scraping is permitted for research purposes and that the process complies with their regulations.
@@
    Follow the university’s rules to submit an ethics approval application.
@@
    Investigate the usage of the APIs offered by the websites. Data scraping can be challenging, especially for websites that store large amounts of data, as searching for user-specified data can be time-consuming. Additionally, different websites may design different APIs with varying parameters (e.g., keywords, time spans) to support efficient data scraping. Therefore, it is important to be familiar with the API usage to ensure that the downloaded data meets expectations.
In machine learning, data preprocessing is a critical step for models to achieve satisfactory performance. This step is also vital for topic modelling because the quality of the original scraped data is often poor and not suitable for direct use in topic models. For example, website links commonly appear in social media posts and news articles; however, they are less informative for topic modelling and may adversely affect model performance. In this research, data preprocessing primarily involves several tasks: stopword removal, punctuation removal, email and URL removal, special symbol removal, lemmatisation, and checking whether the text length exceeds the model’s limit.
Data were collected in JSON format from three primary sources: The Guardian, Reddit, and Twitter. From The Guardian, we extracted full news articles including titles, sections, publication dates, URLs, and complete content. Reddit data comprised post titles, creation timestamps, URLs, and main content (self-text), excluding comments. Twitter data included both original tweets and retweets, encompassing user screen names, creation timestamps, tweet IDs, and full tweet text. Comments and replies on Reddit and Twitter were not included in our dataset to maintain focus on primary content related to circular economy topics.
@@ 3.3 Hyperparameter Optimisation for Topic Modelling
Hyperparameters in machine learning are parameters that must be set before training or fitting a model. In other words, a machine learning algorithm is configured with a specific set of hyperparameter values at the time of instantiation. These values control the behavior of the algorithm and influence key aspects such as model complexity and convergence. Consequently, different sets of hyperparameter values can lead to varying results. Hyperparameter selection is inherently an iterative process of trial and error, and it typically relies on the evaluation of a quantitative metric.
3.3.1 SOO
Single-objective hyperparameter optimisation is a common step in machine learning research, as most tasks in machine learning follow standard evaluation criteria. For example, in classification tasks, accuracy is often prioritised. In our research, to compare different sets of hyperparameter values, we used the coherence ($C_{\text{NPMI}}$) value, which ranges from -1 to 1; a value of 1 indicates the best coherence.
The SOO of topic modeling hyperparameters is defined in Equation 22.
$\boldsymbol{\lambda^{*}}=\underset{\boldsymbol{\lambda}\in\boldsymbol{\Lambda}}{\operatorname{argmin}}\mathbf{V}\left(\mathcal{L},\boldsymbol{\mathcal{A}}_{\boldsymbol{\lambda}},D\right),$ (22)
given a textual dataset $D$, the goal is to find the optimal hyperparameter solution $\boldsymbol{\lambda^{*}}$, where $\mathbf{V}\left(\mathcal{L},\boldsymbol{\mathcal{A}}_{\boldsymbol{\lambda}},D\right)$ measures the quality of the topics generated by the topic modeling approach $\boldsymbol{\mathcal{A}}$ with a hyperparameter solution $\boldsymbol{\lambda}$ on dataset $D$. The quality is determined by a predefined objective function $\mathcal{L}$. $\boldsymbol{\lambda}$ is a vector of length $N$, representing a hyperparameter solution with $N$ hyperparameters (e.g., the number of top words, ngram range), and $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\times\ldots\Lambda_{N}$ is the search space.
Figure 4: Grid Search
To search for the optimal solution in the single-objective case, we propose using grid search in this research. An example of grid search for two hyperparameters is shown in Figure 4. For each combination (represented as a point in the example) of hyperparameters, the model is trained and evaluated using a chosen performance metric. This approach facilitates the identification of the optimal hyperparameter set by comparing the performance across all combinations. The primary advantage of grid search is its simplicity and thoroughness. It ensures that all possible combinations of the specified hyperparameter values are explored, guaranteeing that the optimal set within the grid is found. This exhaustive approach is particularly beneficial when the hyperparameter space is relatively small, as it allows for a comprehensive evaluation of model performance across all parameter settings. In our research, we limited the size of the grid search space to expedite the discovery of the optimal solution, but the search space was primarily designed based on literature review to ensure effectiveness.
3.3.2 MOO
As previously mentioned, SOO has inherent limitations. Unlike other machine learning tasks, evaluating generated topics in topic modelling involves various metrics and approaches, making it a more complex and subjective process. To address this issue, we propose applying multi-objective hyperparameter optimisation (MOO) to the hyperparameters of topic models to provide additional supportive information through another set of topics with different constraints.
MOO is a process used to optimise two or more conflicting objectives simultaneously. Unlike single-objective optimisation, where a single criterion is maximised or minimised, MOO seeks to find a balance among multiple objectives, often resulting in a set of trade-off solutions known as Pareto optimal solutions. Each solution in the Pareto front represents a situation where improving one objective would lead to a deterioration in another.
The MOO of topic modelling hyperparameters is defined in Equation 23.
$P=\{\boldsymbol{\lambda}_{1}^{*},\boldsymbol{\lambda}_{2}^{*},...\}=\underset{\boldsymbol{\lambda}\in\boldsymbol{\Lambda}}{\operatorname{argmin}}\mathbf{V}\left(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{A}_{\boldsymbol{\lambda}},D\right),$ (23)
given a textual dataset $D$, the goal is to find the Pareto set $P$ of optimal solution $\{\boldsymbol{\lambda}_{1}^{*},\boldsymbol{\lambda}_{2}^{*},...\}$, where $\mathbf{V}\left(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{A}_{\boldsymbol{\lambda}},D\right)$ measures the quality of the topics generated by the topic modelling approach $\boldsymbol{\mathcal{A}}$ with a hyperparameter solution $\boldsymbol{\lambda}$ on dataset $D$. The quality is determined by predefined objective functions $\mathcal{L}_{1},\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ which respectively correspond to coherence, diversity, and perplexity in our research. $\boldsymbol{\lambda}$ is a vector of length $N$, representing a hyperparameter solution with $N$ hyperparameters (e.g., the number of top words, ngram range), and $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\times\ldots\Lambda_{N}$ is the search space.
To search the Pareto set, we proposed to use a Tree-structured Parzen Estimator (TPE). TPE [66] is a probabilistic model-based optimisation algorithm often used for hyperparameter tuning. TPE models the search space as two distributions: one for promising hyperparameter values (those expected to perform well) and another for less promising values [67]. It iteratively refines the search by focusing on the promising regions, making TPE more efficient than grid or random search, particularly when dealing with complex search spaces. MOO-TPE [50] is an extension of the TPE algorithm that efficiently handles multiple conflicting objectives. MOO-TPE leverages the concept of Pareto dominance to find better nondomination ranked solutions to form the promising set. However, in MOO, the results is a set of solutions which does not fit our research, as our target is to find one hyperparameter solution. Therefore, in our research, we proposed to use an approach based ideal point [52] concept to find the best hyperparameter solution, the approach is defined as the Equation 24.
$\boldsymbol{\lambda}_{best}=\sum_{n=1}^{|P|}\mathbf{U}(\boldsymbol{\lambda_{i}},\boldsymbol{s(\lambda}^{*}_{n}\boldsymbol{)}).$ (24)
The vector $\boldsymbol{\lambda_{i}}$ represents the ideal point, also known as the ideal solution. For example, in our research, the ideal point is $[1,1,1]$; each dimension corresponds to one objective (e.g., coherence, diversity, and perplexity), and each value represents the best possible value for that objective. $\mathbf{U}$ is the distance function used to quantify the distance between the solutions $\boldsymbol{\lambda}^{*}_{n}$ within the Pareto set $P$ and $\boldsymbol{\lambda_{i}}$. In this research, we used the Euclidean distance, inspired by the work presented in [52]. It is worth noting that the value ranges of different objective functions may vary, so all values need to be rescaled by the function $\boldsymbol{s}$ into the same range before calculating the distances. In our research, $\boldsymbol{s}$ is implemented using MinMaxScaler, scaling each feature to $[0,1]$.